OpenStudy (anonymous):

can anyone help me solve this question PLEASE??? Create a linear equation in two variables in the form y = mx + b to graph using the slope intercept form

OpenStudy (anonymous):

Just plug numbers into the m and b

OpenStudy (anonymous):

ok hun that is easier sad than done never done this and dont know how to do it :(

OpenStudy (anonymous):

Can you tell me two numbers?

OpenStudy (anonymous):

ummmmm just like any two numbers ?

OpenStudy (anonymous):

yeah any!

OpenStudy (anonymous):

Your favorite numbers if you want.

OpenStudy (anonymous):

7 and 2

OpenStudy (anonymous):

Ok, then lets m = 7 and b = 2.

OpenStudy (anonymous):

so, \[y=7x+2\]

OpenStudy (anonymous):

ok now how would I work that ?

OpenStudy (anonymous):

You mean, how do you solve the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You can solve it by graphing or doing algebra.

OpenStudy (anonymous):

Do you know how to graph that equation?

OpenStudy (anonymous):

yes I kind of got how to do that part

OpenStudy (anonymous):

Well I you graph the equation just look at the point where the line crosses the x-axis.

OpenStudy (anonymous):

it crosses the x-axis at some value of x. That "some value of x" is the solution of the equation.

OpenStudy (anonymous):

ok can you help me out with this question? Create a linear equation which requires two steps to solve

OpenStudy (anonymous):

You already have it:

OpenStudy (anonymous):

\[7x+2=0\]

OpenStudy (anonymous):

step 1:\[7x=-2\]

OpenStudy (anonymous):

step 2: \[x = -\frac{2}{7}\]

OpenStudy (anonymous):

ok cool now the last one I need help with is . Create a linear equation in two variables in the form Ax + By = C to graph using the x- and y- intercept method.

OpenStudy (anonymous):

You need to do the same, plug values into the A, B, and C. Can you do that?

OpenStudy (anonymous):

so like this 2a+4b=c ?right

OpenStudy (anonymous):

Not exactly, but you are close. I meant plug values into the A, B, C, for example: \[2x+4y=4\]

OpenStudy (anonymous):

ok now can you walk me through trying to solve that please?

OpenStudy (anonymous):

First, do you know what it means to solve an equation?

OpenStudy (anonymous):

kind of but it is all kind of new to me I never went to high school I did home schooling